OpenStudy (swag):

Write the following value in standard notation. 4.25 x 10^6 425,000 4,250,000 425,610 425,000,000

OpenStudy (swag):

@jazy

OpenStudy (anonymous):

Is it 4.25 x 10^6

OpenStudy (anonymous):

@SWAG

OpenStudy (swag):

Yeah

OpenStudy (anonymous):

\[4.25 \times 10^6=4250000 \]

OpenStudy (anonymous):

10^6 = 10 * 10 * 10 * 10 * 10 * 10 = 1,000,000 1,000,000 * 4.25 = 4,250,000

OpenStudy (swag):

So is it A or B?

OpenStudy (anonymous):

4,250,000

OpenStudy (swag):

SWAG